OpenStudy (anonymous):

If you are finding the line perpendicular to x + 3y = 5, what is the slope of the perpendicular line?

OpenStudy (anonymous):

It will always be the negative reciprocal of the line give. So first put that equation into standard form. That will be y=-1/3x + 5/3 So that slope is -1/3 so perpendicular to that is 3, so 3 is the slope

OpenStudy (anonymous):

The slope is perpendicular to the original slope. To begin, get your equation in slope-intercept form. x + 3y = 5 -> y = -1/3x + 5/3 So you see your slope is -1/3 By the definition of perpendicular, your perpendicular slope is the negative reciprocal of 3. M=3 is the perpendicular slope of the new line.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

You're welcome (: